# Ceresa cycles of bielliptic Picard curves

### v1

@@ Abstract.
We show that the Ceresa cycle $\kappa(C_{t})$ of the genus $3$ curve $C_{t}\colon y^{3}=x^{4}+2tx^{2}+1$ is torsion if and only if $Q_{t}=(\sqrt[3]{t^{2}-1},t)$ is a torsion point on the elliptic curve $y^{2}=x^{3}+1$. This shows that there are infinitely many smooth plane quartic curves over ${\rm\mathbb{C}}$ (resp. ${\rm\mathbb{Q}}$) with torsion (resp. infinite order) Ceresa cycle. Over $\overline{{\rm\mathbb{Q}}}$, we show that the Beilinson–Bloch height of $\kappa(C_{t})$ is proportional to the Neron–Tate height of $Q_{t}$. Thus, the height of $\kappa(C_{t})$ is nondegenerate and satisfies a Northcott property. To prove all this, we show that the Chow motive that controls $\kappa(C_{t})$ is isomorphic to $\mathfrak{h}^{1}$ of an appropriate elliptic curve.
@@ 1. Introduction
Let $C$ be a smooth, projective, geometrically integral curve of genus $g\geq 2$ over an algebraically closed field $k$. We embed $C$ in its Jacobian $J$ using a $(2g-2)$-th root of the canonical bundle. The (canonical) Ceresa cycle of $C$ is the class
$\kappa(C)=[C]-(-1)^{*}[C]\in{\rm CH}_{1}(J)$
in the Chow group of $1$-cycles on $J$ with ${\rm\mathbb{Z}}$-coefficients. [^1] This depends on our choice of $(2g-2)$-th root, but only up to a $(2g-2)$-torsion class; see Lemma 2.1. While $\kappa(C)$ is homologically trivial, Ceresa [12] famously showed that it need not be trivial in ${\rm CH}_{1}(J)$. Indeed, if $k={\rm\mathbb{C}}$ then the Ceresa cycle of a very general genus $g\geq 3$ curve has infinite order. On the other hand, if $C$ is hyperelliptic then $\kappa(C)=0$. Beyond this, very little is known about the set of genus $g$ curves with vanishing or torsion Ceresa cycle, though see §1.4 for some recent results and examples.
In this paper, we study the Ceresa cycles of bielliptic Picard curves over fields of characteristic $\neq 2,3$ [28]. As $k$ is algebraically closed, such a curve has an affine model of the form
@@
for some $t\in k\setminus\{\pm 1\}$, unique up to replacing $t$ with $-t$. Our main result shows that the Ceresa cycles of these curves are governed by points on the single elliptic curve $\widehat{E}\colon y^{2}=x^{3}+1$.
@@
Let $C_{t}$ be the smooth projective curve with model (1.1), and let $Q_{t}$ be the point ($3]{t^{2}-1},t)\in\widehat{E}(k)$, for some choice of cube root. Then $\kappa(C_{t})$ is torsion in ${\rm CH}_{1}(J_{t})$ if and only if $Q_{t}$ is torsion in $\widehat{E}(k)$.
Theorem 1.1 (which is proved in §5.5) appears to be the first characterization of torsion Ceresa cycles in a positive-dimensional family of curves, aside from families where $\kappa(C)$ is identically torsion. We deduce from this result the following corollary (see §5.6).
@@
There are infinitely many ${\rm\overline{\mathbb{Q}}}$-isomorphism classes of genus $3$ curves over ${\rm\mathbb{Q}}$ with infinite order Ceresa cycle. In fact, $\kappa(C_{t})$ has infinite order for all $t\in{\rm\mathbb{Q}}\setminus\{0,\pm 1,\pm 3\}$.
Corollary 1.2 is not surprising given Ceresa’s result, but it seems to have been open until now and is still open in higher genus, as far as we know.
On the other hand, since $\widehat{E}$ has infinitely many torsion points, we also deduce from Theorem 1.1 that there are infinitely many plane quartic curves over ${\rm\mathbb{C}}$ with torsion Ceresa cycle. In fact, our proof shows that there is an algebraic correspondence from $J_{t}$ to $\widehat{E}$ that sends $\kappa(C_{t})$ to a constant multiple of $Q_{t}$. This allows us to deduce the following stronger statement:
@@
There exist infinitely many plane quartic curves over ${\rm\mathbb{C}}$ with torsion Ceresa cycle. Moreover, their orders are unbounded.
Note that it was only recently discovered that there exist any plane quartic curve with torsion Ceresa cycle: Qiu–Zhang [37, §4.1] have shown that $C_{0}\colon y^{3}=x^{4}+1$ has this property.
Interestingly, all known examples in Theorem 1.3 can be defined over ${\rm\overline{\mathbb{Q}}}$; for the curves $C_{t}$, this follows immediately from Theorem 1.1. For $t\in{\rm\overline{\mathbb{Q}}}$, we write $h(\kappa(C_{t}))$ for the Beilinson–Bloch height $\langle\kappa(C_{t}),\kappa(C_{t})\rangle$ [6, 27] of $\kappa(C_{t})$. We show that $h(\kappa(C_{t}))$ is proportional to the Neron–Tate height of $Q_{t}$ (Theorem 5.19), which implies the following nondegeneracy and Northcott properties.
@@
We have $h(\kappa(C_{t}))=0$ if and only if $\kappa(C_{t})$ is torsion. Moreover, for any $X\in{\rm\mathbb{R}}$ and $d\in{\rm\mathbb{N}}$, the set
[…]
@@
Since the field of definition of $C_{t}$ is ${\rm\mathbb{Q}}(t^{2})$ (Lemma 3.2), this implies that for every fixed $d\in\mathbb{N}$, there are finitely many ${\rm\overline{\mathbb{Q}}}$-isomorphism classes of curves $C_{t}$ with torsion Ceresa cycle that can be defined over a number field of degree $\leq d$. In particular, there are finitely many torsion Ceresa cycles $\kappa(C_{t})$ defined over ${\rm\mathbb{Q}}$. In fact, there are three such curves up to ${\rm\overline{\mathbb{Q}}}$-isormorphism: $C_{0}$, $C_{\pm 3}$ and $C_{\pm\sqrt{-3}}$ (Proposition 5.18). Note that $C_{\sqrt{-3}}$ has ${\rm\mathbb{Q}}$-model $y^{3}=x^{4}+6x^{2}-3$.
Finally, we show that for all $t$, the image of $\kappa(C_{t})$ in the Griffiths group of homologically trivial cycles modulo algebraic equivalence is torsion, in the following strong sense:
@@
There exists an integer $N\geq 1$ such that for every bielliptic Picard curve $C$ over a field of characteristic not $2$ nor $3$, the cycle $N\cdot\kappa(C)\in{\rm CH}_{1}(J)$ is algebraically trivial.
@@ 1.1. Relation with modified diagonal cycles
All results above hold verbatim for the canonical modified diagonal cycle $\Delta_{GS}(C_{t})\in{\rm CH}_{1}(C_{t}^{3})$ defined by Gross and Schoen [21]; see §2 for the definition. Indeed, $\kappa(C)$ is torsion if and only if $\Delta_{GS}(C)$ is torsion [44, Theorem 1.5.5], and our method of proof implies a variant of Theorem 1.4 for $\Delta_{GS}(C)$ as well.
S. Zhang has shown a similar Northcott property for the height of $\Delta_{GS}(C)$ in families of smooth curves over smooth projective bases [44, Theorem 1.3.5], i.e. compact subvarieties inside the moduli space $\mathcal{M}_{g}$. As the family $C_{t}$ does not extend to a compact curve inside $\mathcal{M}_{3}$, Theorem 1.4 suggests that Zhang’s Northcott property might very well extend to an open dense subset of $\mathcal{M}_{g}$.
@@ 1.2. Summary of proof
A curve $C$ of the form (1.1) has an order $6$ automorphism (or more canonically, a $\mu_{6}$-action), and the class $\kappa(C)$ is fixed by the induced $\mu_{6}$-action on $J$. On the other hand, over ${\rm\mathbb{C}}$, the $\mu_{6}$-invariant part of the Hodge structure $\operatorname{H}^{3}(J({\rm\mathbb{C}}),{\rm\mathbb{Q}})(1)$ is two-dimensional and of type $(1,0)+(0,1)$. This suggests that there should exist an algebraic correspondence from $J_{t}$ to an elliptic curve that sends $\kappa(C_{t})$ to a point on this elliptic curve. Instead of attempting to construct this correspondence explicitly, we prove this by maneuvering within the flexible category of (pure) Chow motives.
@@
where $J$ is the Jacobian of any curve of the form $C\colon y^{3}=x^{4}+ax^{2}+b$, $E^{\Delta}$ is a certain twist of $\widehat{E}$ depending on $a$ and $b$, and the remainder of the notation will be explained later. It is crucial for our proof that we find a universal such isomorphism, in the category of relative Chow motives over the parameter space $S={\rm\mathbb{A}}^{2}_{{\rm\mathbb{Z}}[1/6]}\setminus\{\Delta_{a,b}=0\}$. The elliptic scheme $\mathcal{E}^{\Delta}\to S$ can then be thought of as the universal $j$-invariant 00 elliptic curve with a nonzero marked point $\mathcal{Q}$. The main ingredients in the construction of this universal isomorphism are the Chow–Kunneth decomposition for motives of abelian schemes (via Beauville’s Fourier transform) [3, 13], Moonen’s refinement of this decomposition in the presence of extra endomorphisms [32], and our study of the Galois action on the endomorphism algebra of the universal Prym surface $\mathcal{P}\to S$ [28]. The latter endomorphism algebra is a nonsplit quaternion algebra over ${\rm\mathbb{Q}}$.
The Ceresa cycle of $C$ naturally lives in ${\rm CH}^{2}(\mathfrak{h}^{3}(J)^{\mu_{6}})$ (Lemma 5.12), so the question becomes: what is the corresponding element of ${\rm CH}^{2}(\mathfrak{h}^{1}(E^{\Delta}(-1))=E^{\Delta}(k)\otimes{\rm\mathbb{Q}}$, under the isomorphism (1.2)? We show that it corresponds to a ${\rm\mathbb{Q}}^{\times}$-multiple of the marked point $\mathcal{Q}$ by showing it universally over $S$, using two soft inputs. First, the elliptic scheme $\mathcal{E}^{\Delta}\rightarrow S$ (which after ${\rm\mathbb{G}}_{m}$-scaling, is essentially an elliptic surface) has Mordell–Weil group of rank 1, generated by $\mathcal{Q}$. Second, we show by explicit computation that the universal Ceresa cycle has at least one infinite order specialization, using reduction mod $p$ arguments based on the methods of [8, 11, 15], see §2. Therefore $\kappa(\mathcal{C}/S)$ corresponds to a nonzero section of $\mathcal{E}^{\Delta}$, and hence a multiple of $\mathcal{Q}$. Specializing at geometric points, we recover Theorem 1.1. The construction allows us to quickly deduce the remaining results in the introduction as well.
@@ 1.3. Geometric interpretation via bigonal duality
$C_{t}\colon y^{3}=x^{4}+2tx^{2}+1$ is a double cover of the elliptic curve $E_{t}\colon y^{3}=x^{2}+2tx+1$. Associated to any such double cover is the bigonal dual cover $\widehat{C}_{t}\to\widehat{E}_{t}$ constructed by Pantazis and Donagi [36, 14], whose Prym is dual to the Prym of $C_{t}\to E_{t}$; see [28, §2.7] for a construction in the situation considered here. In our case, the curve $\widehat{C}_{t}$ has model $y^{3}=(x^{2}+t)^{2}-1$ so that $\widehat{E}_{t}$ is isomorphic to $\widehat{E}\colon y^{2}=x^{3}+1$, for all $t$. The branch locus of $\widehat{C}_{t}\to\widehat{E}$ consists of $\infty$ and the three points ($3]{t^{2}-1},t)$. Theorem 1.1 therefore says that the Ceresa cycle $\kappa(C_{t})$ is torsion if and only if the bigonal dual $\widehat{C}_{t}$ is branched along torsion points of $\widehat{E}$. We find this formulation intriguing, but our proof does not actually use the bigonal dual construction in a direct way.

### v2

@@ Abstract.
We prove that the Chow class $\kappa_{\infty}(C_{t})$ of the Ceresa cycle of the genus three curve $C_{t}\colon y^{3}=x^{4}+2tx^{2}+1$ is torsion if and only if $Q_{t}=(\sqrt[3]{t^{2}-1},t)$ is a torsion point on the elliptic curve $y^{2}=x^{3}+1$. In particular, there are infinitely many plane quartic curves over ${\rm\mathbb{C}}$ with torsion Ceresa cycle. Over $\overline{{\rm\mathbb{Q}}}$, we show that the Beilinson–Bloch height of $\kappa_{\infty}(C_{t})$ is proportional to the Néron–Tate height of $Q_{t}$. Thus, the height of $\kappa_{\infty}(C_{t})$ is nondegenerate and satisfies a Northcott property. To prove all this, we show that the Chow motive that controls $\kappa_{\infty}(C_{t})$ is isomorphic to $\mathfrak{h}^{1}$ of an appropriate elliptic curve.
@@ 1. Introduction
Let $k$ be an algebraically closed field and $C$ a smooth, projective, and connected curve over $k$ of genus $g\geq 2$ with Jacobian variety $J$. Let $e$ be a degree-$1$ divisor on $C$ and let $\iota_{e}\colon C\hookrightarrow J$ be the Abel-Jacobi map based at $e$. We study the torsion behaviour of the Ceresa cycle
(1.1) $\kappa_{e}(C):=[\iota_{e}(C)]-(-1)^{*}[\iota_{e}(C)]\in{\rm CH}_{1}(J)$
in the Chow group modulo rational equivalence. If $\kappa_{e}(C)$ is torsion, then $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)\otimes{\rm\mathbb{Q}}$, where $K_{C}$ is the canonical divisor class. Moreover if $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)\otimes{\rm\mathbb{Q}}$, then the image of $\kappa_{e}(C)$ in ${\rm CH}_{1}(J)\otimes{{\rm\mathbb{Q}}}$ is independent of $e$ and we denote it by $\kappa(C)$; see §2.1 for these claims. The class $\kappa(C)$ vanishes if and only if $\kappa_{e}(C)$ is torsion for some degree-$1$ divisor $e$.
While $\kappa_{e}(C)$ is always homologically trivial, it need not be trivial in ${\rm CH}_{1}(J)$. Indeed, if $k={\rm\mathbb{C}}$ then Ceresa famously showed [12] that for a very general curve of genus $g\geq 3$, $\kappa_{e}(C)$ is of infinite order for every choice of $e$, even modulo algebraic equivalence. On the other hand, if $C$ is hyperelliptic and $e$ is a Weierstrass point, then $\kappa_{e}(C)=0$. Beyond this, little is known about the set of genus-$g$ curves with a vanishing Ceresa cycle, though see §1.4 for some recent results and examples.
In this paper, we study the Ceresa cycles of bielliptic Picard curves over fields of characteristic $\neq 2,3$ [32]. As $k$ is algebraically closed, such a curve $C$ has an affine model of the form
@@
for some $t\in k\setminus\{\pm 1\}$, unique up to replacing $t$ with $-t$. The unique point $\infty\in C(k)$ at infinity has the property that $(2g-2)\infty=4\infty$ is canonical. Our main result shows that the Ceresa cycles of these curves are governed by points on the single elliptic curve $\widehat{E}\colon y^{2}=x^{3}+1$.
@@
Let $C_{t}$ be the smooth projective curve with model (1.2), and let $Q\in\widehat{E}(k)$ be a point with $y$-coordinate equal to $t$. Then $\kappa_{\infty}(C_{t})\in{\rm CH}_{1}(J_{t})$ is torsion if and only if $Q\in\widehat{E}(k)$ is torsion.
In other words, $\kappa_{\infty}(C_{t})$ is torsion if and only if ($3]{t^{2}-1},t)\in\widehat{E}(k)$ is torsion, for any choice of cube root. See Theorem 5.17 for a slightly different formulation of this theorem which also relates the torsion orders. Theorem 1.1 (which is proved in §5.5) appears to be the first characterization of torsion Ceresa cycles in a positive-dimensional family of curves, aside from families where $\kappa(C)$ is identically zero. We deduce from this result the following corollary (see §5.6).
@@
There are infinitely many ${\rm\overline{\mathbb{Q}}}$-isomorphism classes of genus $3$ curves $C$ with $\kappa(C)\neq 0$ that can be defined over ${\rm\mathbb{Q}}$. In fact, $\kappa_{\infty}(C_{t})$ has infinite order for all $t\in{\rm\mathbb{Q}}\setminus\{0,\pm 1,\pm 3\}$.
The first line of Corollary 1.2 is not surprising given Ceresa’s result, but it does not follow from it. Robin de Jong has given a different proof of this infinitude, using a different family of curves [13].
Since $\widehat{E}$ has infinitely many torsion points, we also deduce from Theorem 1.1 that there are infinitely many plane quartic curves over ${\rm\mathbb{C}}$ with torsion Ceresa cycle. As we explain below, our proof shows that there is an algebraic correspondence from $J_{t}$ to $\widehat{E}$ that sends $\kappa(C_{t})$ to a multiple of $Q$. This allows us to deduce the following stronger statement:
@@
There exist infinitely many $t\in{\rm\mathbb{C}}$ such that $\kappa_{\infty}(C_{t})$ is torsion. Each such $t$ lies in ${\rm\overline{\mathbb{Q}}}$. Moreover, the finite torsion orders of the classes $\kappa_{\infty}(C_{t})$ are unbounded.
Note that it was only recently discovered that there exists any plane quartic curve with torsion Ceresa cycle: Qiu–Zhang [42, §4.1] showed that $C_{0}\colon y^{3}=x^{4}+1$ has this property.
For $t\in{\rm\overline{\mathbb{Q}}}\setminus\{\pm 1\}$, we write $h(\kappa(C_{t}))$ for the Beilinson–Bloch height $\langle\kappa(C_{t}),\kappa(C_{t})\rangle$ [6, 31] of $\kappa(C_{t})$. We show that $h(\kappa(C_{t}))$ is proportional to the Néron–Tate height of $Q$ (Theorem 5.20), which implies the following nondegeneracy and Northcott properties.
@@
We have $h(\kappa(C_{t}))=0$ if and only if $\kappa_{\infty}(C_{t})$ is torsion. Moreover, for any $X\in{\rm\mathbb{R}}$ and $d\in{\rm\mathbb{N}}$, the set
$\left\{t\in{\rm\overline{\mathbb{Q}}}\setminus\{\pm 1\}\colon h(\kappa(C_{t}))\leq X\hskip 8.53581pt\mbox{and}\hskip 8.53581pt[{\rm\mathbb{Q}}(t):{\rm\mathbb{Q}}]\leq d\right\}$
@@
Since the field of definition of $C_{t}$ is ${\rm\mathbb{Q}}(t^{2})$ (Lemma 3.2), this implies that for every fixed $d\in\mathbb{N}$, there are finitely many ${\rm\overline{\mathbb{Q}}}$-isomorphism classes of curves $C_{t}$ with torsion Ceresa cycle that can be defined over a number field of degree $\leq d$. In particular, there are finitely many curves $C_{t}$ with $\kappa_{\infty}(C_{t})$ torsion that can be defined over ${\rm\mathbb{Q}}$. In fact, there are three such curves up to ${\rm\overline{\mathbb{Q}}}$-isomorphism: $C_{0}$, $C_{\pm 3}$ and $C_{\pm\sqrt{-3}}$ (Proposition 5.19). Note that $C_{\sqrt{-3}}$ has ${\rm\mathbb{Q}}$-model $y^{3}=x^{4}+6x^{2}-3$.
Finally, we show that for all $t$, the image of $\kappa_{\infty}(C_{t})$ in the Griffiths group of homologically trivial cycles modulo algebraic equivalence is torsion, in the following strong sense:
@@
There exists an integer $N\geq 1$ such that for every bielliptic Picard curve $C$ over a field of characteristic not $2$ nor $3$, the cycle $N\cdot\kappa_{\infty}(C)\in{\rm CH}_{1}(J)$ is algebraically trivial.
@@ 1.1. Relation with modified diagonal cycles
All results above hold verbatim for the canonical modified diagonal cycle $\Delta_{GS,\infty}(C_{t})\in{\rm CH}_{1}(C_{t}^{3})$ defined by Gross and Schoen [24]; see §2 for the definition of $\Delta_{GS,e}(C)$ and the class $\Delta_{GS}(C)\in{\rm CH}^{3}(C)\otimes{\rm\mathbb{Q}}$. Indeed, $\kappa(C)$ vanishes if and only if $\Delta_{GS}(C)$ does [49, Theorem 1.5.5], and the relation $h(\Delta_{GS}(C))=\frac{3}{2}h(\kappa(C))$ [14, (1.1)] gives an analogue of Theorem 1.4 as well.
S. Zhang has shown a similar Northcott property for the height of $\Delta_{GS}(C)$ in families of curves over smooth projective bases [49, Theorem 1.3.5], i.e. compact subvarieties inside $\mathcal{M}_{g}$. As the family $C_{t}$ does not extend to a compact curve inside $\mathcal{M}_{3}$, Theorem 1.4 suggests that Zhang’s Northcott property might very well extend to an open dense subset $U_{g}$ of $\mathcal{M}_{g}$. Indeed, several months after the first version of this paper was posted, Gao and Zhang proved that such a set $U_{g}$ exists [22]. Interestingly, in another preprint we have shown that the locus of Picard curves must be contained in the complement of $U_{3}$ [33, §1.3], so that Theorem 1.4 does not follow from [22].
@@ 1.2. Summary of proof
A curve $C$ of the form (1.2) has an order $6$ automorphism (or more canonically, a $\mu_{6}$-action), and the class $\kappa_{\infty}(C)$ is fixed by the induced $\mu_{6}$-action on $J$. On the other hand, over ${\rm\mathbb{C}}$, the $\mu_{6}$-invariant part of the Hodge structure $\operatorname{H}^{3}(J({\rm\mathbb{C}}),{\rm\mathbb{Q}})(1)$ is two-dimensional and of type $(1,0)+(0,1)$. This suggests that there should exist an algebraic correspondence from $J$ to an elliptic curve that sends $\kappa_{\infty}(C)$ to a point on this elliptic curve. Instead of attempting to construct this correspondence explicitly, we prove this by maneuvering within the flexible category of (pure) Chow motives.
@@
where $J$ is the Jacobian of any curve of the form $C\colon y^{3}=x^{4}+ax^{2}+b$, $E^{\Delta}$ is a certain twist of $\widehat{E}$ depending on $a$ and $b$, and the remainder of the notation will be explained later. It is crucial for our proof that we find a universal such isomorphism, in the category of relative Chow motives over the parameter space $S={\rm\mathbb{A}}^{2}_{{\rm\mathbb{Z}}[1/6]}\setminus\{\Delta_{a,b}=0\}$. The elliptic scheme $\mathcal{E}^{\Delta}\to S$ can then be thought of as the universal $j$-invariant 00 elliptic curve with a nonzero marked point $\mathcal{Q}$. The main ingredients in the construction of this universal isomorphism are the Chow–Künneth decomposition for motives of abelian schemes (via Beauville’s Fourier transform) [3, 15], Moonen’s refinement of this decomposition in the presence of extra endomorphisms [37], and our study of the Galois action on the endomorphism algebra of the universal Prym surface $\mathcal{P}\to S$ [32]. The latter endomorphism algebra is a nonsplit quaternion algebra over ${\rm\mathbb{Q}}$.
The Ceresa cycle of $C$ naturally lives in ${\rm CH}^{2}(\mathfrak{h}^{3}(J)^{\mu_{6}})$ (Lemma 5.12), so the question becomes: what is the corresponding element of ${\rm CH}^{2}(\mathfrak{h}^{1}(E^{\Delta}(-1))=E^{\Delta}(k)\otimes{\rm\mathbb{Q}}$, under the isomorphism (1.3)? We show that it corresponds to a ${\rm\mathbb{Q}}^{\times}$-multiple of the marked point $\mathcal{Q}$ by showing it universally over $S$, using two soft inputs. First, the elliptic scheme $\mathcal{E}^{\Delta}\rightarrow S$ (which after ${\rm\mathbb{G}}_{m}$-scaling, is essentially an elliptic surface) has Mordell–Weil group of rank 1, generated by $\mathcal{Q}$. Second, we show by explicit computation that the universal Ceresa cycle $\kappa_{\infty}(\mathcal{C})$ has at least one infinite order specialization, using reduction mod $p$ arguments based on the methods of [8, 11, 17], see §2. Therefore $\kappa_{\infty}(\mathcal{C})$ corresponds to a nonzero section of $\mathcal{E}^{\Delta}$, and hence a multiple of $\mathcal{Q}$. Specializing at geometric points, we recover Theorem 1.1. The construction allows us to quickly deduce the remaining results in the introduction as well.
@@ 1.3. Geometric interpretation via bigonal duality
$C_{t}\colon y^{3}=x^{4}+2tx^{2}+1$ is a double cover of the elliptic curve $E_{t}\colon y^{3}=x^{2}+2tx+1$. Associated to any such double cover is the bigonal dual cover $\widehat{C}_{t}\to\widehat{E}_{t}$ constructed by Pantazis and Donagi [41, 16], whose Prym is dual to the Prym of $C_{t}\to E_{t}$; see [32, §2.7] for a construction in the situation considered here. In our case, the curve $\widehat{C}_{t}$ has model $y^{3}=(x^{2}+t)^{2}-1$ so that $\widehat{E}_{t}$ is isomorphic to $\widehat{E}\colon y^{2}=x^{3}+1$, for all $t$. The branch locus of $\widehat{C}_{t}\to\widehat{E}_{t}$ consists of $\infty$ and the three points ($3]{t^{2}-1},t)$. Theorem 1.1 therefore says that the Ceresa cycle $\kappa_{\infty}(C_{t})$ is torsion if and only if $\widehat{C}_{t}$ is branched along torsion points of $\widehat{E}_{t}$. While this formulation is intriguing, our proof does not use the bigonal dual construction in a direct way.
One might ask if this is a more general phenomenon: if $C\rightarrow E$ is a double cover of a genus $1$ curve by a genus $3$ curve, whose ramification locus contains a hyperflex point $\infty\in C(k)$, are the torsion properties of $\kappa_{\infty}(C)$ related to the torsion properties of the branch locus of the dual cover $\widehat{C}\rightarrow\widehat{E}$? Adam Logan has verified for us, using [17], that out of a sampling of 99 such bielliptic curves whose dual cover is branched along $7$-torsion points, all of them have infinite order Ceresa cycle. Thus, the answer seems to be ‘no’ in general. It would be interesting to test this hypothesis on other families of genus $3$ curves lying in the intersection (within $\mathcal{M}_{3}$) of the bielliptic locus and a Picard modular surface associated to an imaginary quadratic field other than ${\rm\mathbb{Q}}(\sqrt{-3})$, but these families are non-trivial to write down. In the aforementioned [33], we have generalized Theorem 1.1 to all genus three Picard curves $y^{3}=f(x)$ using different methods. This suggests that the bielliptic structure which we exploit in this paper, while a very useful tool to understand and “compute” the Ceresa cycle, does not tell the whole story.